Tech Staff Summary, 2/11/2000

Current Priorities in addition to routine tasks

GPS
Chris and Jim will observer I-beam emplacement at Piggot and East Prairie and hang boxes etc on Feb 21 and 22. The remainder of the week will be spent making these sites operational.
Portageville upgrade is complete and no problems have been encountered with any currently operating stations.
The prison site north of Pickwick won't work. Chris will scout farther north for potential sites.

Ripley will be nearly impossible without sacrificing perfect sky-view

Digital Backbone
Water District agreement is somewhere between Nashville and Memphis.
Two radios and dishes are in hand. Antenna cables are on backorder.
Mitch has reviewed the technical manuals and the configuration is very similar to the pipeline-50's. Steve or Mitch will attempt to configure and test using the supplied mini-wip antennas.
When dish antenna cables arrive, we will test Richardson to CERI link, and any other links that need testing.
The Arp forest tower is owned by TWRA (TN Wildlife Resource Agency) so we can get a permission if the western path won't work.
Still trying to get in touch with the Osceola tower owner.
For connection between Aironet bridges (previously referred to as radios) and computers there a couple choices. Greg is researching shielded Mondo twisted pair for the ideal 10base-2 to 10base-2. Mitch/Steve/Carl will research adapters to allow 10base-5 to thicknet to 10base-2.
Greg is still waiting on spec for routers (needed at Frechman's Bayou) from Steve. After reading the Tech Manuals, Mitch thinks this can be done with a dumb hub.
Greg will provide Mitch with lat/lon pairs for both western and eastern path options (now that TAP is running again).
Greg ran a couple path profiles for the Memphis Library (owners of Frenchman's Bayou tower). This should help smooth the permission process if the legal muckity-mucks don't get in the way.

Marked Tree Broadbands
GNAR is installed. Jim will replace serial receive cable. Steve is working with Evans at Gosnell to hammer out firewall problems.
HBAR long period noise problem was fixed. Jim and Greg will visit again to investigate HHZ pegged negative problem. Replace if necessary (ain't guralps fun?).

ET Last count, John had 8 permissions in hand and provided lat/lon pair estimates to Mitch.
Mitch will plot up permitted sites and get with Greg to verify telemetry path suitability.
Mitch will schedule a site visit for he and Greg to go to Morristown and Sewanee. If all goes well, then we can plan installation of node at Morristown and CMG3 at Sewanee.
Still waiting on permit to move RICH.
The water leak for the S-13 Horizontals remains a mystery. It has yet to be reproduced. Repairs to BCRT.N, BHT.N, GMG.E, and RCG.E can proceed in hopes it was a fluke. PII causes should be checked before opening vaults.
Six complete 3-c sets of S-13's now calibrated and ready to go. Bryant will try to get another 3 sets ready by March 6. Mitch will arrange with Carl to provide Bryant with alternate means of getting email. Then Bryant's computer can remain dos and will be dedicated to doing nothing but running calibrations.
Jim will make sure sufficient hardware is ready.
Chris will help John install DYTN and take as much hardware and sensors as are ready on March 6.
John is still searching for a helper.

MPH time marks
Still intermittent outages. Probably the amp relays.

NC truck
Chris will get an estimate on truck/winch/shell.

Second tier priorities:

CERI power systems
All seisnet systems are in the racks, tied down, and backed up. We're ready for the big one.
Greg has requested a wiring diagram for the computer systems and Mitch has pledged to provide it by the 32nd day of December in the year 2009.
Greg has one rack left to diagram.
Should test seisnet computer system loads.
Trench to CERI computers has been cleaned out.
Need to examine generator loads/backup power for general CERI systems (ugh, hope this doesn't mess up the seisnets).

Autozone
Dorman's $2800 is in the bank.
Two NI 16-bit PCI cards are on the way from Golden if they work, Mitch/Steve will order PC and mux.
Greg will refurbish one of the old GPS clocks that supplies the wrong time code. While this time code won't work with the DAQ system, we will record it on one channel in the event absolute time is needed. Data will be stamped with the pc time since, under normal circumstances, only relative time is important.
Jim will service the free-field sensor. Then, after the DAQ is up and we get some data, he will service the other sensors as required.

Other Broadbands
CMG-3/DM-24 packages have been running fine in Mitch's office. They are ready for installation.
Potential sites for these sensors are Sewanee and Morristown, TN.
Where the hell is the borehole sensor?
Greg will contact DTA to get DM-16 returned whether or not the "loop" problem has been fixed. We haven't seen it in the field.

Aftershock Array
Greg has specs and prototype and is awaiting an account number to purchase and build the k2 power/etc systems.
MEQ's
The L-4's from NC will replace the 4.5Hz geophones currently being used. Will need converters to do this.
Permission to borrow necessary items from P/PII has been obtained. Any borrowed items should come from Panda rather that PII if possible.

ISIS upgrade
As time permits, stations will be upgraded to the new 4-channel system with emphasis on the backbone stations (this'll make Harley happy to get clipped data).

As time permits priorities.

Panda/PandaII
No deployments are firm. Proposals have been submitted which could result in deployments no sooner than 10/2000. PII on south campus should be inventoried. Then, the 15 PII packages at CERI should be fixed.

SFTN power upgrade
done.

SFTN general upgrade
This would be very nice, but makes Mitch nervous. It requires pulling the sensor.
Jim will get the horizontal channels on the air for recording at CERI.

Analog backbone reconfigured is waiting on the lawyers.

EBZ/PWLA telemetry. Another repeater?

LNXT node
Secure propane tank.
Secure racks.
replace coax jumpers
replace rg-58 unshielded jumpers

MKTA node

NMAD node

Station book. Jim has good records for each station but it would be nice to have an IRIS style station book.

Board Shop
General maintenance.
Fix duct holes.
Reconfigure exhaust that UM took apart?
Move to darkroom area? (this would likely fix non-functional shower problem).

Backburner priorities

Status channels.
Currently supplied on last 8 channels.

LNXT node
permit/install L7 and L8

MKTA node
permit/install remaining 3 ISIS

NMAD node